Sol Summary – February 17th

Sol: 7

Sol Summary title: Moon Cheese and White Mars Bars

Author: Daniel Robson

Mission status: Nominal

Sol activity summary: This morning began in spectacular fashion thanks to the heroic efforts of Hannah, our crew’s Communication Officer, who woke up early to make us all cinnamon buns to start the day right. The morning’s EVA team was geared up for their expedition in good time and left only a few minutes later than scheduled. They drove Opportunity and Curiosity to within a few hundred metres of Copernicus Highway in the Moon Region before pulling over. The rest of their EVA was accomplished on foot in the Grey Moon area, and succeeded in bringing back samples from 9 different sites – including the fossilised shells’ of ancient sea creatures! The EVA team were monitoring the weather and as they noticed heavy clouds approaching (likely with snow) from the southern mountain and skyline ridge, they decided to return to Hab. The reunited crew enjoyed a tasty soup of cheese and broccoli with added chicken for lunch. Not long after we finished, the crew noticed snow flurries beginning from the Hab’s windows. After 10 minutes or so they hadn’t stopped and so the decision was taken to cancel the afternoon’s EVA. Even after the snow eventually stopped, from the looks of the clouds surrounding our little oasis, we all agreed it was probably safer to have stayed and worked in the base itself.

In the afternoon extra measures were taken to secure the maintenance of the base, including repairing any loose tunnel coverings (one had already begun blowing away), fixing the GreenHab door and later the EVA suits that have some faulty fans. For tonight, to keep away the cold, we have some Mars burgers made by Hannah featuring lettuce leaves from our very own GreenHab! After dinner, we’re hopefully going to find a movie from the newly reorganised DVD collection, perhaps even a Christmas film to reflect the weather. The weather outside may be frightful, but inside our home on Mars at MDRS, the company is most definitely delightful…

Look ahead plan: If the weather outside tomorrow is safe, the crew intends to conduct an EVA, deep into Condor Chasma to retrieve samples from the harder to reach areas. The team for this will be mostly made up from those that didn’t get the chance to do an EVA this afternoon. This is provided the snow holds off overnight too. Now that the Journalist (Maria) has compiled all the crew interview videos she is hoping to start editing together our Crew Video, but for today and tomorrow it is her turn to be XO and CO. As always, science will continue in earnest at the base. Many samples from today and likely tomorrow morning are still to be analysed and catalogued fully, and also the final few Arabidopsis seedlings have been transferred to be grown in regolith substrate as part of our botanical ISRU experiment.

Anomalies: None

Weather: cloudy with snow clouds and flurries moving in from the South

Reports to be filed: Sol Summary, EVA report, Operations Report, Greenhab Report, Journalist Report, Astronomy Report, Science Report

Copyright © The Mars Society. All rights reserved. | Main Site